A 0.1500 M HCl solution was prepared on a day when the temperature was 20oC. What is concentration of the solution when used the next day at 27oC? The density of water is 0.9982071 g/mL at 20oC and 0.9965162 g/mL at 27oC.
Answer #1

consider 1 Kg of solution
at 20 oC,
density = 0.9982071 Kg/L
Volume = mass/ density = 1/0.9982071 = 1.00179612 L
number of moles of HCl = M*V
=0.1500 M * 1.00179612 L
= 0.150269418 mol

at 27 oC,
density = 0.9965162 Kg/L
Volume = mass/ density = 1/0.9965162 = 1.003495979 L

M = number of moles/volume
= 0.150269418 /1.003495979
= 0.1497 M
Answer: 0.1497 M